1. In a resonance tube where longitudinal stationery waves can be formed,the first and next sec loud sound is heard when the length of air column is 16cm and 49cm .calc velocity of sound in air ,if frequency of air molecules is 512Hz

2. The absolute zero is _273celcius .what do you understand by this


1
Expert's answer
2021-04-29T10:43:15-0400

1. The velocity of sound in air is


"v=2(L_1-L_2)f=2(0.49-0.16)512=338\text{ m/s}."

2. By absolute zero we understand the temperature at which all mechanical oscillations/motion of particles that compose a body stops.